OpenStudy (anonymous):

Please Help!!!!!!!!!! Ill Give Medal Which rate is a unit rate? A.\[\frac{ 500 Miles }{ 25 Gallons }\] B. \[\frac{ 15 Pounds }{ 3 Cans }\] C. \[\frac{ $12 }{ 1Hour }\] D. \[\frac{ $1 }{ 2Pastries }\]

OpenStudy (anonymous):

Which rate is a unit rate? A.\[\frac{ 500 Miles }{ 25 Gallons }\] B. \[\frac{ 15 Pounds }{ 3 Cans }\] C.\[\frac{ $12 }{ 1Hour }\] D. \[\frac{ $1 }{ 2Pastries }\]

OpenStudy (anonymous):

C

OpenStudy (freckles):

which ever one says 1 on the bottom